OpenStudy (anonymous):

Find the number of right triangles that exist such that each side length is a positive perfect square.

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@terenzreignz @Callisto

terenzreignz (terenzreignz):

Well, suppose we have such a triangle :) Then its dimensions could be neatly represented as \(\large a^2\) , \(\large b^2\) and \(\large c^2\) where a, b, and c are integers :)

terenzreignz (terenzreignz):

Without loss of generality, let's have \(\large c^2\) be the hypotenuse

terenzreignz (terenzreignz):

That can only mean, by the Pythagorean theorem, that \[\Large (a^2)^2 + (b^2)^2 = (c^2)^2\] \[\Large a^4 +b^4 = c^4\] Where a, b, and c are integers...

terenzreignz (terenzreignz):

Something is amiss, don't you think? ^_^

Directrix (directrix):

number of right triangles --> My conjecture is zero exist.

terenzreignz (terenzreignz):

Conjecture? Their existence kind of contradicts Ol' Fermat :D

OpenStudy (anonymous):

Okay, because there are sums of squares that are squares themselves, and there are sums of cubes that are cubes themselves, but nothing higher, right?

terenzreignz (terenzreignz):

sums of cubes that are cubes themselves? Not of two cubes.... three cubes, but never 2. case in point: Fermat's Last Theorem

OpenStudy (anonymous):

OK, great, thanks.

terenzreignz (terenzreignz):

:)